Adigirl29:

The aquarium has 6 fewer yellow fish than green fish. 40 percent of the fish are yellow. How many green fish are in the aquarium? Show your work.

Sherixn:

If 40% are yellow how many are green

Adigirl29:

uh idk

Sherixn:

There's 100% in total

Adigirl29:

uhh its saying to show work

Adigirl29:

it has to do with solving rational equations

Sherixn:

If you don't need my help that's fine

Adigirl29:

im just saying I don't think its that simple-

Sherixn:

I'm trying to guide you through it, that's what we have to do on this website. I just can't give you the answer

Adigirl29:

I don't know what to do sorry

Adigirl29:

I just need help on how to get the answer what do I have to do

Sherixn:

Ok so set up an equation

Sherixn:

yellow fish = green fish - 6

Sherixn:

which can be represented as y=g-6

Adigirl29:

okay

Sherixn:

y = 40% of y + g (which is the total fish)

Sherixn:

y = 0.4(y+g)

Sherixn:

since we're looking for green fish change y for g-6

Sherixn:

g-6 = 0.4(g-6+g)

Sherixn:

g-6 = 0.4(2g-6)

Sherixn:

expand and solve

Sherixn:

I'll leave the rest to you, but let me know if you need help <3

Adigirl29:

so how would I solve the last part of the equation?

Sherixn:

Expand the right hand side using the distributive property

Sherixn:

Then get g onto 1 side

Sherixn:

And then get everything else onto the other

Sherixn:

Then divide both sides by whatever g is being multiplied by

Adigirl29:

uh so lets start with the first step how would I use distributive property on the right side

Sherixn:

0.4(2g-6) multiply 2g by 0.4 and multiply -6 by 0.4

Sherixn:

0.8g - 2.4

Sherixn:

6-g = 0.8g -2.4 is your new equation

Adigirl29:

ohhhh okay okay I got it so that's the right side

Sherixn:

yep (:

Sherixn:

Now subtract 0.8g from both sides

Sherixn:

0.2g = 3.6

Sherixn:

Divide both sides by 0.2

Sherixn:

Your answer is 18 green fish

Adigirl29:

oh my gosh that makes so much sense

Sherixn:

Glad I could help. I'm sorry for being super confusing in the beginning

Sherixn:

Just ignore the first message I sent lol

Adigirl29:

its perfectly fine, do you think you can help me with one more question?